# Boosting Active Defense Persistence: A Two-Stage Defense Framework Combining Interruption and Poisoning Against Deepfake

### v2

@@ Abstract
Active defense strategies have been developed to counter the threat of deepfake technology. However, a primary challenge is their lack of persistence, as their effectiveness is often short-lived. Attackers can bypass these defenses by simply collecting protected samples and retraining their models. This means that static defenses inevitably fail when attackers retrain their models, which severely limits practical use. We argue that an effective defense not only distorts forged content but also blocks the model’s ability to adapt, which occurs when attackers retrain their models on protected images. To achieve this, we propose an innovative Two-Stage Defense Framework (TSDF). Benefiting from the intensity separation mechanism designed in this paper, the framework uses dual-function adversarial perturbations to perform two roles. First, it can directly distort the forged results. Second, it acts as a poisoning vehicle that disrupts the data preparation process essential for an attacker’s retraining pipeline. By poisoning the data source, TSDF aims to prevent the attacker’s model from adapting to the defensive perturbations, thus ensuring the defense remains effective long-term. Comprehensive experiments show that the performance of traditional interruption methods degrades sharply when it is subjected to adversarial retraining. However, our framework shows a strong dual defense capability, which can improve the persistence of active defense. Our code will be available at https://github.com/vpsg-research/TSDF.
Index Terms:
Deepfake, active defense, two-stage defense, intensity separation, face detection interference.
@@ I Introduction
Figure 1: Illustration of an interruption-based defense and its failure to retrain. The lower path demonstrates an effective interruption process, where protected face data leads to a distorted output from the deepfake model. The upper path illustrates a critical flaw in this interruption-only defense. Attackers can bypass the defenses by retraining their model on the protected images. This adaptation makes the model immune to the interruption, ultimately causing the defense to fail.
@@
To counter these threats, researchers have proposed two types of defense strategies: passive detection and active defense. Passive detection aims to identify a forged image after its creation by analyzing a wide array of forged traces, such as GAN fingerprints, compression artifacts, or inconsistent biological signals [3, 4, 5]. However, this method cannot prevent the production of forgery media in advance. Considering the characteristics of social media, forgeries can spread quickly and cause irreversible damage before they are detected and marked [7, 8]. Furthermore, passive detection methods create a continuous arms race, as new generative models are often explicitly trained to eliminate the artifacts that detectors search for, requiring constant updates to detection techniques. This defect of passive response highlights the necessity of adopting a more active strategy. Consequently, active defense has gained increasing attention, as it aims to prevent malicious content from being created or spread in the first place. The core idea is to intervene before or during content generation to disrupt the forgery process itself. Common methods fall into two main categories: interruption and data poisoning. For the former, the goal of interruption is to distort the forged image at the moment of creation. For example, CMUA [15] embeds perturbations into images to make the final output visibly distorted or unusable [14, 39]. TAFIM method [9] prevents the face from being tampered by changing the model output into a pure color image, which can also be used to identify manipulation techniques. For the latter, the goal of data poisoning is to disrupt the deepfake model in the training stage. Some methods, such as Landmark Breaker [19] and Face Poison [22], achieve this by introducing contaminated data to disrupt key pre-processing steps like facial alignment, thus degrading the quality of the trained model [16].
@@
While both interruption and poisoning strategies are effective in their respective stages, they share the critical challenge of lacking persistence. As illustrated in the upper path of Fig. 1, this weakness becomes apparent when an attacker collects a set of protected images and incorporates them into their training dataset. For example, a comparison of the Structural Similarity Index (SSIM) shows that the output from a retrained model is much less distorted than the output from the original model. The core of the problem lies in the pre-processing stage. Conventional interruption perturbations do not interfere with face detectors, meaning that an attacker’s model still receives the protected face as a valid training input. Consequently, from the model’s perspective, these perturbations are not a disruption to be ignored, but simply another data feature to be learned. Given that the model’s objective is to accurately reproduce its training data, it naturally learns to replicate the perturbation patterns as part of the face itself. This adaptation allows the model to become immune to the perturbations, which significantly reduces the defense’s effectiveness and exposes a fundamental flaw in existing static defense strategies [17, 24]. This shows that the defense is effective only upon initial deployment and fails to provide persistent protection against an adaptive attacker.
@@
To overcome this cycle of defense adaptation and build a new paradigm for achieving a more persistent defense, we argue that the defense system should have dual capabilities. That is, it not only needs to interrupt the current forgery content, but also needs to destroy the attacker’s subsequent adaptability. To this end, we propose a Two-Stage Defense Framework. As shown in Fig. 2, the framework’s key insight is the synergistic combination of two strategies in a single mechanism. This combination is efficiently managed by our proposed intensity separation mechanism, which ensures the two defensive functions can operate cooperatively without conflict. In this design, the poisoning mechanism acts as a persistent shield for the interruption function. In the inference stage, the perturbations act as an interruption tool to distort the output [15]. In the training stage, it also serves as a poisoning perturbation to interfere with face localization. This contaminates the dataset used by the attacker for retraining, thus protecting the interruption function from being bypassed [19, 22].
@@ II-A Single-Function Perturbation Methods
Data poisoning is an active defense strategy that interferes with a deepfake model’s learning process by injecting perturbations into the training data. Prior works have explored various approaches that primarily target specific pre-processing steps to contaminate the training data pipeline. Some methods, such as Landmark Breaker [19], focus on a downstream task by disrupting the extraction of facial key points, which aims to degrade the final model’s quality by providing it with poorly aligned facial data for training. However, this approach targets a downstream task that is based on the successful prior face detection. Other methods attack the more basic stage of the pipeline, such as Face Poison [22] and SSOD [23], which directly target the detection process itself. Failure in this initial detection stage can terminate all subsequent processing steps, such as cropping and landmark alignment, preventing the attacker from acquiring usable facial data in the first place.
Interruption perturbations are an active defense strategy designed to distort the output of a deepfake model during the inference stage. These perturbations are often designed to be universal, meaning they are effective across different models and images. For instance, CMUA-Watermark [15] generates a cross-model universal adversarial watermark, which is achieved by optimizing the single perturbation pattern. When this perturbation pattern is added to the image, it can force deepfake models to produce an obvious distortion effect. Similarly, Tang et al. [20] assume that disrupting foundational features is key to a generalizable defense and specifically attacks an ensemble of common and pre-trained feature extractors. By destroying the fundamental semantic representations that most deepfake models rely on, their defense can effectively distort the quality of the final forged image across a wide range of different model architectures. Other notable works like Anti-forged [39] also operate on similar principles, generating robust, perceptual-aware perturbations in the Lab color space to degrade the visual quality of the forged image.
@@ II-B Multifunctional Perturbation Methods
To provide more comprehensive protection than single-function approaches, some studies have begun to explore multifunctional methods that combine different defensive goals. A common direction has been to merge interruption with traceability. For instance, Dual Defense [25] pioneers a dual-effect defense through a single embedded watermark that synergistically provides traceability while working to disrupt face-swapping models. Similarly, Zhu et al. [26] introduce an information-containing perturbation designed to combat facial manipulation systems by embedding an identity message, thus enabling active forensics within the interrupted process. Wu et al. [27] propose a framework to resolve the common conflict between watermarking and detection. Their method fine-tunes a watermark, increasing a deepfake detector’s accuracy on watermarked images without compromising traceability.
@@
Figure 3: The overall framework of the TSDF method. The process begins in the interruption stage (upper left), where an initial perturbation ($W$) is generated by feeding a perturbed image into several deepfake models and optimizing $W$ to maximize the feature-level Mean Squared Error loss ($L_{MSE}$) and a feature enhancement loss ($L_{enh}$). Subsequently, the poisoning stage (lower left) generates a specialized perturbation to attack multiple face detectors by maximizing feature-level loss ($L_{feat}$) and output-level loss ($L_{output}$). In the combination and optimization stage (right side), the two perturbations are efficiently fused. The combined system uses a threshold $\tau$ to identify low-intensity regions in the interruption perturbations, creating a mask where the poisoning perturbation is selectively applied.
@@ III-A Deepfake Interruption via Perturbation
Instead of directly attacking end-to-end deepfake models like CMUA [15], the proposed method targets feature extractors used in deepfake networks. This prevents feature extractors from correctly encoding facial attributes, thereby disrupting downstream attribute editing and identity reconstruction. Given an input image x, an interruption perturbation W is introduced to generate a perturbed image $\hat{X}$ = x+W. The goal is to maximize the discrepancy between extracted feature representations.
To achieve this, we optimize the interruption perturbation to maximize the distance between extracted features from the original and perturbed images:
@@
$\mathcal{L}_{\mathrm{MSE}}=MSE\sum_{i=1}^{N}(E_{i}(\hat{X}),E_{i}(x)).$ (2)
This loss ensures that the perturbation creates a consistent and significant distortion across the feature extractors of all targeted models. To achieve a more potent disruption than a simple MSE loss can provide, we introduce a feature enhancement loss $\mathcal{L}_{\mathrm{enh}}$ that amplifies discrepancies across three key levels: local, global and structural. This amplification is driven by a non-linear exponential function, which is critical for magnifying even small feature differences into a significant and destructive effect.
$\mathcal{F}_{\mathrm{enh}}=E_{i}(\hat{X})+\alpha\cdot e^{\frac{\|\hat{W}\circ\Delta\|}{\sigma}}\cdot\sum_{i=1}^{3}w_{i}\Delta_{i}.$ (3)
Here, $\mathcal{F}_{\mathrm{enh}}$ denotes the enhanced feature. The term $w_{i}$ is a predefined weight vector that balances these three types of discrepancy. This amplification is regulated by a temperature scaling parameter $\sigma$. The term $\Delta$ represents a vector of local, global, and structural differences $(\mathrm{L_{D}},\mathrm{G_{D}},\mathrm{S_{D}})$, while $\hat{W}$ denotes the adaptive weights that control the enhancement strength. This component ensures that small feature differences are amplified into significant disruptions, resulting in a more comprehensive and destructive feature attack.
@@
$\mathcal{L}_{\mathrm{enh}}=MSE\sum_{i=1}^{N}(\mathcal{F}_{\mathrm{enh}},E_{i}(x)).$ (8)
The feature enhancement loss $\mathcal{L}_{\mathrm{enh}}$ is then formulated to quantify the magnitude of this amplification. It calculates the MSE between the enhanced feature $\mathcal{F}_{\mathrm{enh}}$ and the clean feature $E_{i}(x)$. Optimizing to maximize this loss drives the perturbation to create larger and more diverse feature discrepancies, thus achieving a more powerful disruptive effect.
@@
$\mathcal{L}_{\mathrm{feature}}=\lambda\mathcal{L}_{\mathrm{enh}}+\mathcal{L}_{\text{MSE}}.$ (9)
To maximize the feature loss, we iteratively optimize the perturbation W using gradient ascent. This optimization follows the Projected Gradient Descent (PGD) framework [28], where a projection step is used after each update to ensure the perturbation remains within the imperceptibility budget $\epsilon$:
$W_{t+1}=\mathrm{Proj}_{\epsilon}(W_{t}+\gamma\cdot\nabla_{W}\mathcal{L}_{\mathrm{feature}}),$ (10)
where $\gamma$ is the step size and $\epsilon$ is the perturbation budget. The projection operator ensures that the interruption perturbation remains within a bounded perturbation space.
@@ III-B Poisoning the Retraining Pipeline
To ensure a generalizable poisoning effect against a range of detectors, we employ a joint attack strategy at the feature-level. This principle of attacking across the feature layers of multiple networks has also been explored in methods like Face Poison [22]. As formulated in Equations 11, our approach creates a fused feature representation by combining outputs from the feature layers of multiple SOTA detectors. The feature-level loss $L_{feat}$ is designed to maximize the difference between features within the masked poisoning regions ($M_{d}$). For a batch of images, the fusion features is:
@@
$L_{\text{feat}}=\sum_{i=1}^{N}w_{i}\cdot\|F_{\text{fused}}^{i}-(-F_{\text{fused}}^{i,\text{clean}}+\nu\cdot M_{d})\|^{2}.$ (12)
Here, $D_{i}(\text{x})$ is the feature extracted from the face detector $i_{th}$. $\alpha_{i}$ is the fusion weights that balance the contribution of each of the detectors and $w_{i}$ is the weight of the layer $i_{th}$. $N$ is the number of feature layers to attack. $\nu$ is a hyperparameter that controls the magnitude of the targeted feature-space margin. $F_{\text{fused}}^{i,\text{clean}}$ denotes the fused feature representation derived from the original image. The optimization of this poisoning loss is limited to specific regions $M_{d}$ defined by our intensity separation mechanism, and its technical details are given in the following section (III-C).
@@
$L_{total}=\mathcal{L}_{feat}-\mathcal{L}_{output}.$ (14)
The poisoning perturbation is then optimized to maximize this total loss $L_{total}$, which effectively disrupts face detection. The optimization remains constrained to the masked regions to maintain image quality.

### v3

@@ Abstract
Active defense strategies have been developed to counter the threat of deepfake technology. However, a primary challenge is their lack of persistence, as their effectiveness is often short-lived. Attackers can bypass these defenses by simply collecting protected samples and retraining their models. This means that static defenses inevitably fail when attackers retrain their models, which severely limits practical use. We argue that an effective defense not only distorts forged content but also blocks the model’s ability to adapt, which occurs when attackers retrain their models on protected images. To achieve this, we propose an innovative Two-Stage Defense Framework (TSDF). Benefiting from the intensity separation mechanism designed in this paper, the framework uses dual-function adversarial perturbations to perform two roles. First, it can directly distort the forged results. Second, it acts as a poisoning vehicle that disrupts the data preparation process essential for an attacker’s retraining pipeline. By poisoning the data source, TSDF aims to prevent the attacker’s model from adapting to the defensive perturbations, thus ensuring the defense remains effective long-term. Comprehensive experiments show that the performance of traditional interruption methods degrades sharply when these methods are subjected to adversarial retraining. However, our framework shows a strong dual defense capability, which can improve the persistence of active defense. Our code will be available at https://github.com/vpsg-research/TSDF.
@@ I Introduction
Figure 1: Illustration of an interruption-based defense and its failure to retrain. The lower path demonstrates an effective interruption process, where protected face data leads to a distorted output from the deepfake model. The upper path illustrates a critical flaw in this interruption-only defense. Attackers can bypass the defense by retraining their model on the protected images. This adaptation makes the model immune to the interruption, ultimately causing the defense to fail.
@@
To counter these threats, researchers have proposed two types of defense strategies: passive detection and active defense. Passive detection aims to identify a forged image after its creation by analyzing a wide array of forged traces, such as GAN fingerprints, compression artifacts, or inconsistent biological signals [3, 4, 5, 6, 7, 8, 9]. However, this method cannot prevent the production of forgery media in advance. Considering the characteristics of social media, forgeries can spread quickly and cause irreversible damage before they are detected and marked [10, 11]. Furthermore, passive detection methods create a continuous arms race, as new generative models are often explicitly trained to eliminate the artifacts that detectors search for, requiring constant updates to detection techniques. This defect of passive response highlights the necessity of adopting a more active strategy. Consequently, active defense has gained increasing attention, as it aims to prevent malicious content from being created or spread in the first place [12, 13, 14, 15, 16]. The core idea is to intervene before or during content generation to disrupt the forgery process itself. Common methods fall into two main categories: interruption and data poisoning. For the former, the goal of interruption is to distort the forged images at the moment of creation [18, 17]. For example, CMUA [19] embeds perturbations into images to make the final output visibly distorted or unusable. TAFIM method [20] prevents the face from being tampered by changing the model output into a pure color image, which can also be used to identify manipulation techniques. For the latter, the goal of data poisoning is to disrupt the deepfake model in the training stage. Some methods, such as Landmark Breaker [21] and Face Poison [22], achieve this by introducing contaminated data to disrupt key pre-processing steps like facial alignment, thus degrading the quality of the trained model [23].
@@
While the interruption strategies are effective in the inference stage, they share the critical challenge of lacking persistence. As illustrated in the upper path of Fig. 1, this weakness becomes apparent when an attacker collects a set of protected images and incorporates them into their training dataset. For example, a comparison of the Structural Similarity Index Measure(SSIM) shows that the output from a retrained model is much less distorted than the output from the original model. The core of the problem lies in the pre-processing stage. Conventional interruption perturbations do not interfere with face detectors, meaning that an attacker’s model still receives the protected face as a valid training input. Consequently, from the model’s perspective, these perturbations are not a disruption to be ignored, but simply another data feature to be learned. Given that the model’s objective is to accurately reproduce its training data, it naturally learns to replicate the perturbation patterns as part of the face itself. This adaptation allows the model to become immune to the perturbations, which significantly reduces the defense’s effectiveness and exposes a fundamental flaw in existing static defense strategies [24, 25]. This shows that the defense is effective only upon initial deployment and fails to provide persistent protection against an adaptive attacker.
@@
To overcome this cycle of defense adaptation and build a new paradigm for achieving a more persistent defense, we argue that the defense system should have dual capabilities. That is, it not only needs to interrupt the current forgery content, but also needs to destroy the attacker’s subsequent adaptability. To this end, we propose a Two-Stage Defense Framework. As shown in Fig. 2, the framework’s key insight is the synergistic combination of two strategies in a single mechanism. This combination is efficiently managed by our proposed intensity separation mechanism, which ensures the two defensive functions can operate cooperatively without conflict. In this design, the poisoning mechanism acts as a persistent shield for the interruption function. In the inference stage, the perturbations act as an interruption tool to distort the output [19]. In the training stage, it also serves as a poisoning perturbation to interfere with face localization [22]. This contaminates the dataset used by the attacker for retraining, thus protecting the interruption function from being bypassed.
@@ II-A Single-Function Perturbation Methods
Data poisoning is an active defense strategy that interferes with a deepfake model’s learning process by injecting perturbations into the training data. Prior works have explored various approaches that primarily target specific pre-processing steps to contaminate the training data pipeline. Some methods, such as Landmark Breaker [21], focus on a upstream task by disrupting the extraction of facial key points, which aims to degrade the final model’s quality by providing it with poorly aligned facial data for training. However, this approach targets a upstream task that is based on the successful prior face detection. Other methods attack the more basic stage of the pipeline, such as Face Poison [22] and SSOD [26]. These methods primarily target the face detection and localization process, which serves as the foundational step in most deepfake retraining pipelines. By introducing adversarial perturbations specifically designed to disrupt the classification and bounding box regression of detectors, they can force the detection stage to fail or produce highly inaccurate coordinates. Such failures effectively terminate subsequent pre-processing steps like cropping and landmark alignment, thereby preventing attackers from acquiring usable facial data for model fine-tuning.
Interruption perturbations are an active defense strategy designed to distort the output of a deepfake model during the inference stage. These perturbations are often designed to be universal, meaning they are effective across different models and images. For instance, CMUA-Watermark [19] generates a cross-model universal adversarial watermark, which is achieved by optimizing the single perturbation pattern. When this perturbation pattern is added to the image, it can force deepfake models to produce an obvious distortion effect. Similarly, Tang et al. [27] assume that disrupting foundational features is key to a generalizable defense and specifically attacks an ensemble of common and pre-trained feature extractors. By destroying the fundamental semantic representations that most deepfake models rely on, their defense can effectively distort the quality of the final forged image across a wide range of different model architectures.
@@ II-B Multifunctional Perturbation Methods
To provide more comprehensive protection than single-function approaches, some studies have begun to explore multifunctional methods that combine different defensive goals. A common direction has been to merge interruption with traceability. For instance, Dual Defense [28] pioneers a dual-effect defense through a single embedded watermark that synergistically provides traceability while working to disrupt face-swapping models. Similarly, Zhu et al. [29] introduce an information-containing perturbation designed to combat facial manipulation systems by embedding an identity message, thus enabling traceability within the interrupted process. Wu et al. [30] propose a framework to resolve the common conflict between watermarking and detection. Their method fine-tunes a watermark, increasing a deepfake detector’s accuracy on watermarked images without compromising traceability.
@@
Figure 3: The overall framework of the TSDF method. The process begins in the interruption stage (upper left), where an initial perturbation (W) is generated by feeding a perturbed image into several deepfake models and optimizing W to maximize the feature-level Mean Squared Error loss ($\mathcal{L}_{MSE}$) and minimizing a feature enhancement loss ($\mathcal{L}_{enh}$). Subsequently, the poisoning stage (lower left) generates a specialized perturbation to attack multiple face detectors by maximizing feature-level loss ($\mathcal{L}_{feat}$) and output-level loss ($\mathcal{L}_{output}$). In the combination and optimization stage (right side), the two perturbations are efficiently fused. The combined system uses a threshold $\tau$ to identify low-intensity regions in the interruption perturbations, creating a mask where the poisoning perturbation is selectively applied.
@@ III-A Deepfake Interruption via Perturbation
Instead of directly attacking end-to-end deepfake models like CMUA [19], the proposed method targets feature extractors used in deepfake networks. This prevents feature extractors from correctly encoding facial attributes, thereby disrupting downstream attribute editing and identity reconstruction. Given an input image x, an initial perturbation W is introduced to generate a perturbed image $\hat{X}$ = x+W. The goal is to maximize the discrepancy between extracted feature representations.
To achieve this, we optimize the initial perturbation to maximize the distance between extracted features from the original and perturbed images:
@@
$\mathcal{L}_{MSE}=\frac{1}{N}\sum_{i=1}^{N}\|E_{i}(\hat{X})-E_{i}(x)\|_{2}.$ (2)
This loss ensures that the perturbation creates a consistent and significant distortion across the feature extractors of all targeted models. To achieve a more potent disruption than a simple MSE loss can provide, we introduce a feature enhancement loss $\mathcal{L}_{enh}$ that amplifies discrepancies across three key levels: local, global, and structural. This amplification is driven by a non-linear exponential function, which is critical for magnifying even small feature differences into a significant and destructive effect.
$\mathcal{F}_{enh}=E_{i}(\hat{X})+\alpha\cdot e^{\frac{\|\hat{W}\circ\Delta\|}{\sigma}}\cdot\sum_{i=1}^{3}w_{i}\Delta_{i}.$ (3)
[…]
@@
[…]
$\mathcal{L}_{enh}=\frac{1}{N}\sum_{i=1}^{N}\|\mathcal{F}_{enh}-E_{i}(\hat{X})\|_{2}.$ (8)
By minimizing this specific loss, we force the perturbation to evolve towards larger and more diverse feature discrepancies than a standard MSE loss would achieve, ultimately resulting in a more destructive and generalizable interruption effect.
@@
$\mathcal{L}_{\mathrm{feature}}=\lambda\mathcal{L}_{enh}+\mathcal{L}_{MSE}.$ (9)
To maximize the feature loss, we iteratively optimize the perturbation W using gradient ascent. This optimization follows the Momentum Iterative Fast Gradient Sign Method (MI-FGSM) [31], Unlike standard methods, MI-FGSM integrates a momentum term to stabilize the update direction. The optimization step is formulated as:
$W_{t+1}=\mathrm{Proj}_{\epsilon}(W_{t}+\gamma\cdot\mathrm{sign}(g_{t+1})),$ (10)
where $\gamma$ is the step size and $\epsilon$ is the perturbation budget. The projection operator ensures that the interruption perturbation remains within a bounded perturbation space, and $g_{t+1}$ denotes the accumulated gradient with momentum factor $\mu$.
@@ III-B Poisoning the Retraining Pipeline
To ensure a generalizable poisoning effect against a range of detectors, we employ a joint attack strategy at the feature-level. This principle of attacking across the feature layers of multiple networks has also been explored in methods like Face Poison [22]. As formulated in Equation 11, our approach creates a fused feature representation by combining outputs from the feature layers of multiple state-of-the-art (SOTA) detectors. The feature-level loss $\mathcal{L}_{feat}$ is designed to maximize the difference between features within the masked poisoning regions ($M_{d}$). For a batch of images, the fusion features are:
@@
$\mathcal{L}_{feat}=\sum_{i=1}^{N}w_{i}\cdot\|F_{\text{fused}}^{i}-(-F_{\text{fused}}^{i,\text{clean}}+\nu\cdot M_{d})\|^{2},$ (12)
where $D_{i}(\cdot)$ is the feature extracted from the face detector $i_{th}$. $\alpha_{i}$ is the fusion weights that balance the contribution of each of the detectors and $w_{i}$ is the weight of the layer $i_{th}$. $N$ is the number of feature layers to attack. $\nu$ is a hyperparameter that controls the magnitude of the targeted feature-space margin. $F_{\text{fused}}^{i,\text{clean}}$ denotes the fused feature representation derived from the original image. The optimization of this poisoning loss is limited to specific regions $M_{d}$ defined by our intensity separation mechanism, and its technical details are given in the following section (III-C).
@@
$\mathcal{L}_{total}=\mathcal{L}_{feat}-\mathcal{L}_{output}.$ (14)
The $\mathcal{L}_{total}$ constructs a multi-objective optimization that simultaneously disrupts feature-layer representations ($\mathcal{L}_{feat}$) and detector-level outputs ($\mathcal{L}_{output}$). This dual-level approach constrains both feature consistency and bounding box geometry, thereby increasing the difficulty of model adaptation during retraining. The poisoning perturbation is then optimized to maximize this total loss $\mathcal{L}_{total}$, which effectively disrupts face detection. The optimization remains constrained to the masked regions to maintain image quality.